For a population with σ = 10, a score that is located 20 points above the mean would have a z-score of?
Answer by jim_thompson5910(35256)

z = (x-mu)/sigma

z = ((20+mu)-mu)/10

z = (20+mu-mu)/10

z = (20)/10

z = 2